Shift report 20011117 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 17 2001 - 08:23:02 EST

  • Next message: Flemming Videbaek: "run status"

    ------------------------------------------------------------
    Date: Sat Nov 17 08:23:02 2001, Shift: 00:00-08:00
    Supervisor: Y.K.Lee, with: 
    Log book#11 pages 2 - 4
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam for almost 6 hours although MCR could not steer it for us
    very well.
    
    2. Through out the night FS was at 12 degrees with 2/3A current.
        Run 5847 MRS@ 40 degrees with 4A............181K of treigger-6.
        Run 5851 MRS@ 40 degrees with 4A............102 K
        
        Run 5852 MRS@25 degrees with 4A........quickly aborted because
           its deadtime was 95%.
    
        Run 5853 MRS@90 degrees with 6B..............125 K
    
    3. TPM1 has a bad pad (row6, pad 16). It shows as a red column in
    TPC3d.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 17 2001 - 08:23:38 EST